# Length independent generalization bounds for deep SSM architectures with stability constraints

### v1

@@ A State-Space Model (SSM) is a discrete-time linear dynamical system of the form
In particular, a sufficient (but not necessary) condition for stability is that $A$ is a contraction, i.e. $\|A\|_{2}<1$.
A stable SSM $\Sigma$ is not only robust to perturbations, but it is well-known Antoulas (2005) that its input-output map $\mathcal{S}_{\Sigma}$ acts as a linear operator $\ell^{p,p}(\mathbb{R}^{n_{u}})\to\ell^{r,r}(\mathbb{R}^{n_{y}})$, for any choice $p,r\in\{\infty,2\}$, as for any $\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})$, $\mathcal{S}_{\Sigma}(\mathbf{u})\in\ell^{r,r}(\mathbb{R}^{n_{y}})$. In particular, $\mathcal{S}_{\Sigma}$ has a well-defined induced norm as a linear operator, defined in the usual way, $\left\lVert\Sigma\right\rVert_{p,r}=\sup_{\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})}\frac{\left\lVert\mathcal{S}_{\Sigma}(\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}}{\left\lVert\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}}$. In this paper, we will use the induced norms $\left\lVert\Sigma\right\rVert_{2,\infty}$ and $\left\lVert\Sigma\right\rVert_{\infty,\infty}$, which can be upper bounded by the following two norms defined on SSMs.
Definition 2 (Chellaboina et al. (1999))
@@
Lemma 3 (Chellaboina et al. (1999))
For a system of form (1) $\left\lVert\Sigma\right\rVert_{\infty}\leq\left\lVert\Sigma\right\rVert_{1}$ and $\left\lVert\Sigma\right\rVert_{2,\infty}\leq\left\lVert\Sigma\right\rVert_{2}$.
The norms defined above will play a cruicial role in in the main result of the paper, as they will allow us to bound the Rademacher complexity of the deep SSM model.
Remark 4 (Computing $\left\lVert\Sigma\right\rVert_{i}$,$i=1,2$)
@@
Relationship with continuous-time models. In the literature, SSMs are often derived by discretizing a continuous-time linear differential equation in time (e.g. Gu and Dao (2023) and references therein). If the discretization step $\Delta$ is a fixed constant, then we obtain time-invariant linear system of the form (1). In some models, the discretization step can depend on the input and state of the continuous-time state-space representation. In the latter case, one still obtains a linear discrete-time state-space representation, but then the matrices describing the system equation depend on the current input and state, and hence the matrices are time-dependent. In this paper we will stick to time-invariant models (1), for the sake of simplicity.
Deep SSM models. In this paper, we focus our attention to deep SSM models comoposed of multiple linear SSMs. The main reason is that most of the state-of-the-art architectures for long-range sequences are based on linear system. A key characteristics of deep SSM models are the combination of SSM layers with nonlinear transformations, typically some kind of neural network that is constant in time. The general definition we use is the following.
Definition 5
A DT-SSM block (or simply SSM block) is a function $f^{\text{DTB}}:\ell^{p,q}(\mathbb{R}^{n_{u}})\to\ell^{r,s}(\mathbb{R}^{n_{u}})$ that is composed of a stable SSM followed by a nonlinear transformation that is constant in time. That is, $f^{\text{DTB}}(\mathbf{u})[k]=g(\mathcal{S}_{\Sigma}(\mathbf{u})[k])+\alpha\mathbf{u}[k]$ for some $g:\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{u}}$ for all $k\in\left[T\right]$. The function $g$ is represented as either an MLP or a GLU network (see Definitions 7 and 8).
We incorporate $\alpha$ so that the definition covers residual connections (typically $\alpha$ is either 1 or 0). The previous definition is inspired by the series of popular architectures mentioned in the introduction. Similarly, the definition of deep SSM models is also based on these architectures. A deep SSM model consists of SSM blocks along with an encoder, and a decoder transformation preceded by a time-pooling layer. In practice, it is common to apply some normalization techniques, such as batch or layer normalization. As they are not essential for our result, we omit them from the definition. Indeed, once training is finished, a normalization layer corresponds to applying a neural network with linear activation function, i.e., it can be integrated into one of the neural network layer. Since the objective of PAC bounds is to bound the generalization error for already trained models, for the purposes of PAC bounds normalization layers can be viewed as an additional neural network layer.
Definition 6
A discrete time deep SSM model for classification is a function $f:\ell^{p,q}(\mathbb{R}^{n_{\text{in}}})\to\mathbb{R}^{n_{\text{out}}}$ of the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$. The functions $f^{\text{Enc}}$ and $f^{\text{Dec}}$ are linear transformations which are constant in time, while $f^{\text{B}_{i}}$ is a DT-SSM block for all $i$. By pooling we mean the operation $f^{\text{Pool}}(\mathbf{u})=\frac{1}{T}\sum_{k=1}^{T}\mathbf{u}[k]$, an avarage pooling over the time axis.
For regression tasks, we omit the pooling layer and define the Decoder as a linear transformation that is identical for all time steps.
@@
S4 Gu et al. (2021) LTI, $A=\Lambda-PQ^{*}$ block-diagonal SSM + nonlinear activation
S4D Gu et al. (2022) LTI, $A=-\text{exp}(A_{Re})+i\cdot A_{Im}$ block-diagonal SSM + nonlinear activation
S5 Smith et al. (2022) LTI, diagonal $A$ SSM + nonlinear activation
LRU Orvieto et al. (2023) LTI, diagonal $A$ complex exponential parametrization SSM + MLP skip connection
Table 1: Summary of popular deep SSM models.
@@
Definition 8 (GLU layer Smith et al. (2022))
A GLU layer is a function of the form $GLU:\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})\to\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$ parametrized by a linear operaetor $W$ such that $GLU(\mathbf{u})[k]=GELU(\mathbf{u}[k])\odot\sigma(W(GELU(\mathbf{u}[k])))$, where $\sigma$ is the sigmoid function and GELU is the Gaussian Error Linear Unit Hendrycks and Gimpel (2016).
@@
3.2 Learning problem
We consider the usual supervised learning framework for sequential input data. The considered models, parametrized by $\theta$, are of the form $f_{\theta}:\mathcal{U}^{+}\to\mathcal{Y}$, where $\mathcal{U}\subseteq\mathbb{R}^{n_{\text{in}}}$ is the input space, $\mathcal{U}^{+}$ denotes the set of all finite sequences and $\mathcal{Y}$ is the output space, either a finite set for classification or $\mathcal{Y}\subseteq\mathbb{R}^{n_{\text{out}}}$ for regression. Without loss of generality we assume $n_{\text{out}}=1$. In practice, $\theta$ is usually obtained by some learning algorithm, such as Gradient Flow or (Stochastic) Gradient Descent. In this paper, we are agnostic regarding the origin of $\theta$ as we prove a generalization bound that holds for a set of models, therefore we omit the subscript $\theta$ from the notation of $f$.
For training we will use input sequence of the same length $T$. More precisely, let us denote by $\mathcal{U}^{T}$ the set of all sequences of elements of $\mathcal{U}$ of length $T$. A dataset is an i.i.d sample of the form $S=\{(\mathbf{u}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ where $\mathbf{u}_{i}\in\mathcal{U}^{T}$ and $\mathbf{y}_{i}\in\mathcal{Y}$ for all $i$, where the random sample is with respect to some probability measure $\mathcal{D}$ defined on the $\sigma$-algebra generated by the Borel sets of $\mathcal{U}^{T}\times\mathcal{Y}$, where $\mathcal{U}^{T}\times\mathcal{Y}$ is considered with respect to the standard topology as a subset of $(\mathbb{R}^{n_{\text{in}}})^{T}\times\mathbb{R}^{n_{\text{out}}}$.
An elementwise loss function is of the form $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ and is assumed to be $K_{\ell}$-Lipschitz-continuous, i.e. $|\ell(y_{1},y_{1}^{\prime})-\ell(y_{2},y_{2}^{\prime})|\leq K_{\ell}(|y_{1}-y_{2}|+|y_{1}^{\prime}-y_{2}^{\prime}|)$ for all $y_{1},y_{2},y_{1}^{\prime},y_{2}^{\prime}\in\mathbb{R}$, as well as $\ell(y,y)=0$ for all $y\in\mathbb{R}$. Let $\mathcal{L}_{emp}^{S}(f)=\frac{1}{N}\sum_{i=1}^{N}l(f(\mathbf{u}_{i}),y_{i})$ denote the empirical loss of a model $f$ w.r.t a dataset $S$. We denote the true error by $\mathcal{L}(f)=\mathbf{E}_{(\mathbf{u},y)\sim\mathcal{D}}[\ell(f(\mathbf{u}),y)]$. The generalization error or gap of a model $f$ is defined as $|\mathcal{L}_{emp}^{S}(f)-\mathcal{L}(f)|$. In practice, we can approximate the generalization gap by the empirical generalization gap, i.e. the loss difference on the training data and some test data (see Devroye et al. (2013)).
3.3 Assumptions
Before moving forward to discuss the main result, we summarize the assumptions we make in the paper for the sake of readibility. Hereinafter we denote by $\mathcal{F}$ a set of deep SSM models represented by its direct product of its layerwise parameters. Furthermore, let $\mathcal{E}$ denote the set of all SSM models $\Sigma$ for which there is a model $f\in\mathcal{F}$ such that $\Sigma$ is an SSM layer of $f$.
Assumption 9
@@ 3.3 Assumptions
    Bounded input. There exist $K_{\mathbf{u}}>0$ and $K_{y}>0$ such that for all input trajectory $\mathbf{u}$ we have $\left\lVert\mathbf{u}\right\rVert_{\ell^{2,2}(\mathbf{R}^{n_{\text{in}}})}\leq K_{\mathbf{u}}$ and for all labels we have $|y|\leq K_{y}$
@@
    Bounded Encoder and Decoder. We assume the Encoder and Decoder have bounded operator norms, i.e. $\sup\limits_{W^{\text{Enc}}\in\mathcal{W}_{\text{Enc}}}\left\lVert W^{\text{Enc}}\right\rVert_{2,2}<K_{\text{Enc}}$ and $\sup\limits_{W^{\text{Dec}}\in\mathcal{W}_{\text{Dec}}}\left\lVert W^{\text{Dec}}\right\rVert_{\infty,\beta}<K_{\text{Dec}}$ for $\beta\in\mathbb{N}\cup\{\infty\}$.
Assumption 1. is not restrictive as we consider classification.
@@
The Lipschitzness in Assumption 2. holds for most of the loss functions used in practice. We mention that even the square-loss is Lipschitz on a bounded domain. From the practical aspect, the upped boundedness is also mild, as parameters along the learning algorithm’s trajectory usually make $l$ bounded. In the worst case, $l$ is bounded on a bounded domain due to being Lipschitz.
@@
Assumption 3. is yet again standard in the literature. Even in practical applications the input is usually normalized or standardized as a preprocessing step before learning.
@@
Assumption 4. is the most important one as it plays a central role in our work. The motivation behind this assumption is twofold. First, in practical implementation of SSM based architectures, it is very common to apply some structured parametrization of the matrices of the systems, which leads to learning stable matrices. In many cases, the underlying intention is numerical stability of the learning algorithm, however we argue that the major advantage of such parametrizations is to ensure a stable behavior of the system. Second, similar stability assumptions are standard in control theory.

### v2

@@ Appendix A Stability of SSMs
In particular, a sufficient (but not necessary) condition for stability is that $A$ is a contraction, i.e. $\|A\|_{2}<1$. A stable SSM $\Sigma$ is not only robust to perturbations, but its input-output map can be extended to act on the Banach spaces of infinite sequences, generated by $\ell^{p,q}_{T}(\mathbb{R}^{n_{in}})$, $T\geq 0$.
More precisely, denote by $\ell^{2,2}(\mathbb{R}^{n})$ and $\ell^{\infty,\infty}(\mathbb{R}^{n})$ the Banach spaces generated by the all inifinite sequences over $\mathbb{R}^{n}$ such that the quantities $\left\lVert\mathbf{u}\right\rVert^{2}_{\ell^{2,2}(\mathbb{R}^{n})}=\sum_{k=0}^{\infty}\left\lVert\mathbf{u}[k]\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n})}=\sup\limits_{k}\left\lVert\mathbf{u}[k]\right\rVert_{\infty}$ are well defined and finite. If $\mathbf{u}=\mathbf{u}[0]\ldots,\mathbf{u}[T-1]$ is a finite sequence of length $T$, then we can interpret it as an infinite sequence $\mathbf{u}=\mathbf{u}[0]\ldots,\mathbf{u}[T-1],0,0,\ldots$; elements of which are zero after the $T$th element. With this identification, $\ell_{T}^{p,q}(\mathbb{R}^{n})$ is a close subspace of $\subseteq\ell^{p,q}(\mathbb{R}^{n})$, and $\ell^{p,q}(\mathbb{R}^{n})$ contains no proper closed subspace containing $\bigcup_{T\geq 0}\ell_{T}^{p,q}(\mathbb{R}^{n})$, i.e., $\bigcup_{T\geq 0}\ell_{T}^{p,q}(\mathbb{R}^{n})$ generates $\ell^{p,q}(\mathbb{R}^{n})$.
A stable SSM $\Sigma$ is not only robust to perturbations, but its input-output map can be extended to a linear operator $\mathcal{S}_{\Sigma}:\ell^{p,p}(\mathbb{R}^{n_{u}})\to\ell^{r,r}(\mathbb{R}^{n_{y}})$, for any choice $p,r\in\{\infty,2\}$. More precisely, define $\mathcal{S}_{\Sigma}(\mathbf{u})[T-1]=\mathcal{S}_{\Sigma,T}(\mathbf{u}[0]\cdots\mathbf{u}[T-1])[T-1]$ for all $T>0$. It then follows that for any $\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})$, $\mathcal{S}_{\Sigma}(\mathbf{u})\in\ell^{r,r}(\mathbb{R}^{n_{y}})$, $p,r=2,\infty$, see Antoulas (2005). In particular, $\mathcal{S}_{\Sigma}$ has a well-defined induced norm as a linear operator, defined in the usual way, $\left\lVert\mathcal{S}_{\Sigma}\right\rVert_{p,r}=\sup_{\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})}\frac{\left\lVert\mathcal{S}_{\Sigma}(\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}}{\left\lVert\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}}$. In the sequel, by a slight abuse of notation, we will denote by $\left\lVert\Sigma\right\rVert_{p,r}$ the induced norm $\left\lVert\mathcal{S}\right\rVert_{p,r}$.
As it was mentioned above
$\left\lVert\Sigma\right\rVert_{p,r}=\sup_{T\geq 0}\left\lVert\mathcal{S}_{\Sigma,T}\right\rVert_{p,r}$
Indeed, for any $\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})$, $\left\lVert\mathcal{S}_{\Sigma}(\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}=\lim_{T\rightarrow\infty}\left\lVert\mathcal{S}_{\Sigma,T}(\mathbf{u}[0]\cdots\mathbf{u}[T-1])\right\rVert_{\ell^{r,r}_{T}(\mathbb{R}^{n}_{y})}$, and $\left\lVert\mathbf{u}\right\rVert_{\ell^{p,p}(\mathbb{R}^{n_{u}})}=\lim_{T\rightarrow\infty}\left\lVert\mathbf{u}\right\rVert_{\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})}$ and hence $\sup_{T>0}\left\lVert\mathcal{S}_{\Sigma,T}\right\rVert_{p,r}\geq\left\lVert\Sigma\right\rVert_{p,r}$. Moreover, for any $\mathbf{u}\in\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})$, $\left\lVert\mathcal{S}_{\Sigma}(\mathbf{u})\right\rVert_{\ell^{r,r}(\mathbb{R}^{n_{y}})}\geq\left\lVert\mathcal{S}_{\Sigma,T}(\mathbf{u}[0]\cdots\mathbf{u}[T-1])\right\rVert_{\ell^{r,r}_{T}(\mathbb{R}^{n}_{y})}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{p,p}(\mathbb{R}^{n_{u}})}=\left\lVert\mathbf{u}\right\rVert_{\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})}$, hence $\left\lVert\Sigma\right\rVert_{p,r}\geq\left\lVert\mathcal{S}_{\Sigma,T}\right\rVert_{p,r}$.
In this paper, we will use the induced norms $\left\lVert\Sigma\right\rVert_{2,\infty}$ and $\left\lVert\Sigma\right\rVert_{\infty,\infty}$ to upper bound the Rademacher complexity. In turn, these norms can be upper bounded by the following two standard control-theoretical norms defined on SSMs.
Definition A.2 (Chellaboina et al. (1999)).
@@
Lemma A.3 (Chellaboina et al. (1999)).
For a system of form (1) $\left\lVert\Sigma\right\rVert_{\infty,\infty}\leq\left\lVert\Sigma\right\rVert_{1}$ and $\left\lVert\Sigma\right\rVert_{2,\infty}\leq\left\lVert\Sigma\right\rVert_{2}$.
The norms defined above will play a crucial role in in the main result of the paper, as they will allow us to bound the Rademacher complexity of the deep SSM model.
Remark A.4 (Computing $\left\lVert\Sigma\right\rVert_{i}$,$i=1,2$).
@@
Appendix B Elements of deep SSM models
As it was mentioned above, the principal components of deep SSM models are SSM blocks. Various SSM models used in the literature differ from each other in the way the SSM components are parametrized and in the choice of the non-linear component of SSM blocks, see Table 1 for a summary.
@@ Appendix B Elements of deep SSM models
S4 Gu et al. (2021) $A=\Lambda-PQ^{*}$ block-diagonal SSM + nonlinear activation
S4D Gu et al. (2022) $A=-\text{exp}(A_{Re})+i\cdot A_{Im}$ block-diagonal SSM + nonlinear activation
S5 Smith et al. (2022) diagonal $A$ SSM + nonlinear activation
LRU Orvieto et al. (2023) diagonal $A$ exponential parametrization SSM + MLP skip connection
Table 1: Summary of some popular deep SSM models.
Note that in some papers, the matrices of the SSM component are allowed to be complex valued, but such linear dynamical systems can be replaced by linear dynamical systems defined using real matrices, by doubling the dimension of the state-space and that of the input and the output space.
Remark B.1 (Stability assumptions in the SSM literature).
In some of the cited papers, the discrete-time SSM components were obtained by discretizing internally stable continuous-time linear time-invariant dynamical model in time, using a fixed discretization time step. It is well-known in control theory that the thus obtained discrete-time linear systems of the form (1) are also internally stable. In this way, the majority of literature considers deep SSM models for which the SSM components are internally stable, at least as far as the parametrization used for learning is concerned, as they are internally stable in continuous-time.
@@
Definition B.3 (GLU layer Smith et al. (2022)).
A GLU layer is a function of the form $GLU:\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})\to\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$ parametrized by a linear operator $W$ such that $GLU(\mathbf{u})[k]=GELU(\mathbf{u}[k])\odot\sigma(W(GELU(\mathbf{u}[k])))$, where $\sigma$ is the sigmoid function and GELU is the Gaussian Error Linear Unit Hendrycks and Gimpel (2016).
@@
Appendix C Assumptions
Hereinafter we denote by $\mathcal{F}$ a set of deep SSM models represented by its direct product of its layerwise parameters. Furthermore, let $\mathcal{E}$ denote the set of all SSM models $\Sigma$ for which there is a model $f\in\mathcal{F}$ such that $\Sigma$ is an SSM layer of $f$. First, we restate our assumptions:
Assumption C.1
@@ Appendix C Assumptions
    Bounded input. There exist $K_{\mathbf{u}}>0$ and $K_{y}>0$ such that for any input trajectory $\mathbf{u}$ and label $y$ sampled from $\mathcal{D}$, with probability $1$ we have that $\left\lVert\mathbf{u}\right\rVert_{\ell_{T}^{2,2}(\mathbf{R}^{n_{\text{in}}})}\leq K_{\mathbf{u}}$ and $|y|\leq K_{y}$.
@@
    Bounded Encoder and Decoder. We assume the Encoder and Decoder have bounded operator norms, i.e. $\sup\limits_{W\in\mathcal{W}_{\text{Enc}}}\left\lVert W\right\rVert_{2,2}<K_{\text{Enc}}$ and $\sup\limits_{W\in\mathcal{W}_{\text{Dec}}}\left\lVert W\right\rVert_{\infty,\beta}<K_{\text{Dec}}$ for $\beta\in\mathbb{N}\cup\{\infty\}$.
  6. 6.
    Bounded MLPs and GLUs. We assume that any considered MLP or GLU layer has a bounded parameters, i.e. for an $L$-layer model defined in Definition B.2 we have $\max\limits_{1\leq i\leq L+1}\sup\limits_{W\in\mathcal{W}_{i}}\left\lVert W\right\rVert_{\infty,\infty}<K_{W}$ and $\max\limits_{1\leq i\leq L+1}\sup\limits_{\mathbf{b}\in\mathcal{B}_{i}}\left\lVert\mathbf{b}\right\rVert_{\infty}<K_{\mathbf{b}}$. Moreover, for any GLU layer defined in Definition B.3 we have $\sup\limits_{W}\in\mathcal{W}_{\text{GLU}}\left\lVert W\right\rVert_{\infty,\infty}<K_{\text{GLU}}$.
Assumption 1: is not restrictive as we consider classification.
@@
Assumption 2: The Lipschitzness holds for most of the loss functions used in practice. We mention that even the square-loss is Lipschitz on a bounded domain. From the practical aspect, the upped boundedness is also mild, as parameters along the learning algorithm’s trajectory usually make $l$ bounded. In the worst case, $l$ is bounded on a bounded domain due to being Lipschitz.
@@
Assumption 3: is yet again standard in the literature. Even in practical applications the input is usually normalized or standardized as a preprocessing step before learning.
@@
Assumption 4: is the most important one as it plays a central role in our work. The motivation behind this assumption is twofold. First, in practical implementation of SSM based architectures, it is very common to apply some structured parametrization of the matrices of the systems, which leads to learning stable matrices. In many cases, the underlying intention is numerical stability of the learning algorithm, however we argue that the major advantage of such parametrizations is to ensure a stable behavior of the system. Second, similar stability assumptions are standard in control theory.